Entropy Question-Thermodynamics

100 kJ of heat flows through a window. The inside of the window is 20 deg Celsius. The outside of the window is -10 deg Celsius. What is delta Sgen as a result of this process
Answer #1

The change in entropy of a system due to heat flow is given by

AS

Here, we have to consider the two systems independently.

  • For the inside of the window,

T = 20 degrees Celsius = 20+273 = 293 K

Here, the heat is lost. So, Q is negative.

Q = -100 kJ

So, the change in entropy of the system is

ASin = = -100 * 10/293 = -341.2969J/K

  • For the outside of the window,

T = -10 degrees Celsius = -10+273 = 263 K

Here, the heat is gained. So, Q is positive

Q = 100 kJ

So, the change in entropy of the system is

ASout = = 100 * 10/263 = 380.228J/K

So, the total change in entropy is

Astot = ASin + AS out = = -341.2969 +380.228 = 38.9311J/K